# Limit-case admissibility for positive infinite-dimensional systems

### v1

@@ 2.1 Transferring properties between ordered Banach spaces
Let $X$ and $Z$ be ordered Banach spaces and let $J\in\mathcal{L}(X,Z)$ be bipositive. The following assertions hold.
@@
    The map $J$ is injective.
@@
    If $J(X_{+})$ is dense in $Z_{+}$, then $J^{\prime}$ is bipositive.
@@
    If $Z_{+}$ is generating in $Z$ and $J(X)$ is majorizing in $Z$, then $J^{\prime}(Z^{\prime}_{+})=X^{\prime}_{+}$
Proof 2.2.
(a) Let $x\in X$ and $Jx=0$. Then $Jx\geq 0$ and $J(-x)\geq 0$, so the bipositivity of $J$ implies that $x\geq 0$ and $-x\geq 0$. Hence $x=0$ since $X_{+}\cap-X_{+}=\{0\}$.
(b) Since $J$ is positive, so is $J^{\prime}$. Now let $z^{\prime}\in Z^{\prime}$ and assume that $J^{\prime}z^{\prime}\geq 0$. For every $x\in X_{+}$ this implies that $\langle z^{\prime},Jx\rangle\geq 0$. Since $J(X_{+})$ is by assumption dense in $Z_{+}$, we conclude that $\langle z^{\prime},z\rangle\geq 0$ for all $z\in Z_{+}$, so $z^{\prime}\in Z^{\prime}_{+}$.
(c) Since $J^{\prime}$ is positive, we only need to show the inclusion $J^{\prime}(Z^{\prime}_{+})\supseteq X^{\prime}_{+}$, so let $x^{\prime}\in X^{\prime}_{+}$. The linear mapping $\varphi:J(X)\to\mathbb{R}$, $J(x)\mapsto\langle x^{\prime},x\rangle$ is well-defined and positive by the injectivity – known from (a) – and bipositivity of $J$. Since $J(X)$ is by assumption majorizing in $Z$, the Kantorovich extension theorem [4, Theorem 1.36] implies $\varphi$ can be extended to a positive linear functional $z^{\prime}:Z\to\mathbb{R}$. As $Z_{+}$ is generating, it follows that $z^{\prime}$ is automatically continuous [1, Corollary 6], i.e., $z^{\prime}\in Z^{\prime}_{+}$. Moreover, the equality $\langle z^{\prime},Jx\rangle=\varphi(Jx)=\langle x^{\prime},x\rangle$ for each $x\in X$ implies that $x^{\prime}=J^{\prime}z^{\prime}\in J^{\prime}(Z^{\prime}_{+})$.
We note in passing that the proof of (b) only used the positivity rather than the bipositivity of $J$. The following classical example is simple but illustrates Proposition 2.1 quite well:
@@
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded domain, let $k\in\mathbb{N}$, and let $\mathrm{C}^{k}(\overline{\Omega})$ denote the space of all $k$-times continuously differentiable functions on $\Omega$ whose derivative up to order $k$ extend continuously to $\overline{\Omega}$. This is an ordered Banach space when endowed with the pointwise order and the usual norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\mathrm{C}^{k}}$-norm. We show that its dual cone is (more precisely, can be identified with) the set of all positive finite Borel measures on $\overline{\Omega}$.
Indeed, the canonical embedding $J:\mathrm{C}^{k}(\overline{\Omega})\to\mathrm{C}(\overline{\Omega})$ is bipositive and has majorizing range because it contains the constant function $\operatorname{{\mathbbm{1}}}$. Moreover, employing the density of $\mathrm{C}^{k}(\overline{\Omega})$ in $\mathrm{C}(\overline{\Omega})$ (which is true by the Stone-Weierstraß approximation theorem) and by using shifts by small multiples of the constant function $\operatorname{{\mathbbm{1}}}$, we even have that the cone $\mathrm{C}^{k}(\overline{\Omega})_{+}$ is dense in the cone $\mathrm{C}(\overline{\Omega})_{+}$.
So the dual operator $J^{\prime}$ is bipositive and injective according to Proposition 2.1. Hence, $J^{\prime}$ is a bipositive embedding of the dual space $\mathrm{C}(\overline{\Omega})^{\prime}$ (which can be identified with the space of finite Borel measures on $\overline{\Omega}$) into the dual space of $\mathrm{C}^{k}(\overline{\Omega})$. Hence, Proposition 2.1(c) shows that the cone $\mathrm{C}^{k}(\overline{\Omega})$ consists of (to be precise, can be identified via $J^{\prime}$ with) the set of all positive finite Borel measures on $\overline{\Omega}$. In other words, there are not more positive functionals on $\mathrm{C}^{k}(\overline{\Omega})$ than on $\mathrm{C}(\overline{\Omega})$ since the former space is majorizing in the latter.
In our next result, we show that the first two properties in Table 1 are preserved by the inverse of a positive bounded bijection.
@@
    If the cone of $X$ is the face of its bidual cone, then the same is true for the cone of $Z$.
@@
Proof 2.5.
(a) Suppose that every increasing norm-bounded net in $X_{+}$ is norm-convergent and let $(z_{\alpha})$ be an increasing norm-bounded net in $Z_{+}$. Then the net $(Tz_{\alpha})$ is increasing and norm-bounded in $X_{+}$, and hence it is norm-convergent, i.e., there exists $x\in X$ such that $Tz_{\alpha}\to x$ in $X$. As $T$ is a continuous linear bijection, it follows by the bounded inverse theorem that $z_{\alpha}\to T^{-1}x$ in $Z$, as desired.
(b) Let $k:X\to X^{\prime\prime}$ and $\widetilde{k}:Z\to Z^{\prime\prime}$ be the canonical embeddings. Let $z\in Z_{+}$ and $z^{\prime\prime}\in Z^{\prime\prime}_{+}$ such that $0\leq z^{\prime\prime}\leq\widetilde{k}(z)$. We need to show that $z^{\prime\prime}\in\widetilde{k}(Z)$. Since $T$ is positive, so is the double dual $T^{\prime\prime}:Z^{\prime\prime}\to X^{\prime\prime}$. Therefore, $0\leq T^{\prime\prime}z^{\prime\prime}\leq T^{\prime\prime}\widetilde{k}(z)=k(Tz)$. We may now infer from the assumption, $k(X_{+})$ is a face of $X^{\prime\prime}_{+}$, that $T^{\prime\prime}z^{\prime\prime}\in k(X)$. It follows that $z^{\prime\prime}\in(T^{\prime\prime})^{-1}k(X)=\widetilde{k}(T^{-1}X)=\widetilde{k}(Z)$; here we’ve implicitly used that bijectivity of $T$ implies bijectivity of $T^{\prime\prime}$.
(c) First of all, recall that normality of the cone in an ordered Banach space is equivalent to every order interval being norm-bounded [4, Theorem 2.40]. Now, let $a,b\in Z$. Then $T[a,b]\subseteq[Ta,Tb]$ due to the positivity of $T$ and the later set is norm-bounded because of the normality of the cone in $X$. Hence, $[a,b]=T^{-1}T[a,b]$ is also norm-bounded since $T^{-1}$ is continuous.
For a bipositive operator $J:X\to Z$ between ordered Banach spaces, we are interested in conditions that ensure that $J(X_{+})$ is a face of $Z_{+}$. For this purpose, we introduce the following result regarding extension of linear functionals:
@@
Let $X$ be an ordered Banach space with a generating cone and let $V$ be a vector subspace of $X$. Let $x^{\prime}\in X^{\prime}_{+}$ and $\varphi:V\to\mathbb{R}$ be linear. The following are equivalent.
@@
“(ii) $\Rightarrow$ (i)”: Let (ii) be true. First of all, note that the functional $p:X\to[0,\infty)$ given by $x\mapsto\inf\{\left\langle x^{\prime}\,,\,w\right\rangle:w\in X_{+},w\geq x\}$ is well-defined since $X_{+}$ is generating and can easily checked to be sublinear. Moreover, one has $\varphi(v)\leq p(v)$ for all $v\in V$ since, if $v\in V$ is given, then $\varphi(v)\leq\left\langle x^{\prime}\,,\,w\right\rangle$ for each $w\in X_{+}$ with $w\geq x$, by assumption. We may thus employ the Hahn-Banach theorem to obtain a linear functional $y^{\prime}:X\to\mathbb{R}$ that satisfies $y^{\prime}\raisebox{-2.15277pt}{$|$}_{V}=\varphi$ and $y^{\prime}(x)\leq p(x)$ for all $x\in X$. For each $x\in X_{+}$, observe that $y^{\prime}(-x)\leq p(-x)=0,$ so $y^{\prime}$ is positive. In particular, $y^{\prime}\in X^{\prime}$ since the cone $X_{+}$ is generating [1, Corollary 6]. Finally,
$\left\langle y^{\prime}\,,\,x\right\rangle\leq p(x)\leq\left\langle x^{\prime}\,,\,x\right\rangle,\quad(x\in X_{+})$
proves that $y^{\prime}\leq x^{\prime}$.
@@
Let $X$ be an ordered Banach space with a generating and normal cone such that $X_{+}$ is a face of $X^{\prime\prime}_{+}$. Also, let $Z$ be an ordered Banach space such that there exists a bipositive embedding $J:X\to Z$.
Assume that $(R_{n})$ is a sequence in $\mathcal{L}(Z,X)_{+}$ such that $(JR_{n})$ and $(R_{n}J)$ both converge to the identity operator (on $Z$ and $X$ respectively) in the weak operator topology. Then $J(X_{+})$ is a face of $Z_{+}$.
Proof 2.9.
Denoting $k:X\to X^{\prime\prime}$ and $\widetilde{k}:Z\to Z^{\prime\prime}$ by the canonical embeddings, the following diagram commutes:
@@
Moreover, density of $J$ implies the injectivity of $J^{\prime}:Z^{\prime}\to X^{\prime}$. Let $x\in X_{+}$ and $z\in Z_{+}$ such that $0\leq z\leq Jx$. Define $\varphi:J^{\prime}(Z^{\prime})\to\mathbb{R}$ as
$v\mapsto\varphi(v):=\left\langle\widetilde{k}(z)\,,\,(J^{\prime})^{-1}v\right\rangle.$
Let $v=J^{\prime}z^{\prime}\in J^{\prime}(Z^{\prime})$ and fix $w\in X^{\prime}_{+}$ such that $w\geq v$. Then
$\varphi(v)=\left\langle\widetilde{k}(z)\,,\,z^{\prime}\right\rangle=\left\langle z^{\prime}\,,\,z\right\rangle=\lim_{n\to\infty}\left\langle z^{\prime}\,,\,JR_{n}z\right\rangle=\lim_{n\to\infty}\left\langle v\,,\,R_{n}z\right\rangle.$
Positivity of $z$ and $w$ thus implies that
$\varphi(v)\leq\limsup_{n\to\infty}\left\langle w\,,\,R_{n}z\right\rangle\leq\limsup_{n\to\infty}\left\langle w\,,\,R_{n}Jx\right\rangle=\left\langle w\,,\,x\right\rangle=\left\langle k(x)\,,\,w\right\rangle.$
Thus by Proposition 2.6, there exists $y^{\prime\prime}\in X^{\prime\prime}$ such that $0\leq y^{\prime\prime}\leq k(x)$ and $y^{\prime}\raisebox{-2.15277pt}{$|$}_{J^{\prime}(Z^{\prime})}=\varphi$. Furthermore, for each $z^{\prime}\in Z^{\prime}$, we obtain that
$\left\langle\widetilde{k}(z)\,,\,z^{\prime}\right\rangle=\varphi(J^{\prime}z^{\prime})=\left\langle y^{\prime\prime}\,,\,J^{\prime}z^{\prime}\right\rangle=\left\langle J^{\prime\prime}y^{\prime\prime}\,,\,z^{\prime}\right\rangle$
and consequently, $\widetilde{k}(z)=J^{\prime\prime}y^{\prime\prime}$.
Finally, since $k(X_{+})$ is a face of $X^{\prime\prime}_{+}$ and $0\leq y^{\prime\prime}\leq k(x)$, there exists $y\in X$ such that $y^{\prime\prime}=k(y)$. As $k$ is bipositive, it follows that $0\leq y\leq x$. Wherefore for each $z^{\prime}\in Z^{\prime}$, we have
$\left\langle z^{\prime}\,,\,z\right\rangle=\left\langle\widetilde{k}(z)\,,\,z^{\prime}\right\rangle=\left\langle J^{\prime\prime}k(y)\,,\,z^{\prime}\right\rangle=\left\langle\widetilde{k}(Jy)\,,\,z^{\prime}\right\rangle=\left\langle z^{\prime}\,,\,Jy\right\rangle$
which proves that $z=Jy\in J(X)$, as desired.
We show in Example 2.17 that the assumption $X_{+}$ is a face of $X^{\prime\prime}_{+}$ cannot be dropped in Theorem 2.8.
If $X$ and $D$ are Banach spaces, it is a well-known consequence of the Hahn-Banach extension theorem that every operator $T\in\mathcal{L}(X,D)$ satisfies the norm equality $\left\lVert T^{\prime}\right\rVert_{D^{\prime}\to X^{\prime}}=\left\lVert T\right\rVert_{X\to D}$. The following theorem can be interpreted as a variation of the inequality $\leq$” in this equality. For an intuition about the theorem, we refer to the subsequent Examples 2.1.
@@
for a number $c\geq 0$ that might depend on all involved spaces and operators except for $X$ and $T$. Here, $\operatorname{span}(E^{\prime}_{+})=E^{\prime}_{+}-E^{\prime}_{+}$ is endowed with the norm given in formula (1.2) (with $X_{+}$ replaced with $E^{\prime}_{+}$).
Note that the space $D$ does not occur in the conclusion of the theorem. We are interested in the mappings $\tilde{J}T:X\to\tilde{E}$ and $(JT)^{\prime}:E^{\prime}\to X^{\prime}$ and $D$ is an auxiliary space to relate the properties of those two mappings. Before giving the proof of the theorem, we illustrate two situations where the assumptions of the theorem are fulfilled:
{examples}
Let $p\in[1,\infty]$, let $T:L^{p}(\mathbb{R})\to L^{\infty}(\mathbb{R})$ be a positive bounded linear operator, and assume that every function in the range of $T$ is Lipschitz continuous. Let $\operatorname{Lip}(\mathbb{R})$ denote the space of all scalar-valued Lipschitz continuous functions on $\mathbb{R}$, which is an ordered Banach space when endowed with the pointwise order and the norm $\left\lVert f\right\rVert_{\operatorname{Lip}}:=\left\lvert f(0)\right\rvert+\operatorname{Lip}(f)$, where $\operatorname{Lip}(f)$ is the Lipschitz constant of $f$. Then Theorem 2.10 implies that
$\left\lVert T^{\prime}\right\rVert_{\operatorname{span}(\operatorname{Lip}(\mathbb{R})^{\prime}_{+})\to(L^{p})^{\prime}}\leq c\left\lVert T\right\rVert_{L^{p}\to L^{\infty}}$
for a constant $c$ that does not depend on $T$.
To see this, choose $X=L^{p}(\mathbb{R})$, $E=\operatorname{Lip}(\mathbb{R})$, $\tilde{E}=L^{\infty}(\mathbb{R})$ and $D=E\cap\tilde{E}$ with the norm $\left\lVert f\right\rVert_{D}:=\left\lVert f\right\rVert_{\operatorname{Lip}}+\left\lVert f\right\rVert_{\infty}$ in the theorem, let $J$ and $\tilde{J}$ be the canonical embeddings, and observe that $T$ is bounded from $L^{p}(\mathbb{R})$ to $D$ by the closed graph theorem.

### v4

@@ 2.1. Transferring properties between (pre-)ordered Banach spaces
Let $J:X\to Z$ be a bounded operator between pre-ordered Banach spaces $X$ and $Z$.
@@
    If $X$ is an ordered Banach space and $J$ is bipositive, then $J$ is injective.
@@
    The map $J^{\prime}$ is bipositive if and only if $J(X_{+})$ is a dense subset of $Z_{+}$.
@@
    If $Z_{+}$ is generating in $Z$ and $J$ is bipositive, injective, and has majorizing range, then $J^{\prime}(Z^{\prime}_{+})=X^{\prime}_{+}$
Proof.
(a) Let $x\in X$ such that $Jx=0$. Then $Jx\geq 0$ and $J(-x)\geq 0$, so the bipositivity of $J$ implies that $x\geq 0$ and $-x\geq 0$. Hence, $x=0$ since $X_{+}\cap-X_{+}=\{0\}$.
(b) Let $J^{\prime}$ be bipositive. As $J^{\prime}$ is positive, so is $J$ and hence, $J(X_{+})\subseteq Z_{+}$. Assume that the inclusion is not dense. By the Hahn-Banach separation theorem, there exists $z\in Z_{+}$ and $z^{\prime}\in Z^{\prime}$ such that $\left\langle z^{\prime}\,,\,z\right\rangle<\left\langle z^{\prime}\,,\,Jx\right\rangle=\left\langle J^{\prime}z^{\prime}\,,\,x\right\rangle$ for all $x\in X_{+}$. Taking $x=0$ we obtain $\left\langle z^{\prime}\,,\,z\right\rangle<0$, so $z^{\prime}\not\in Z^{\prime}_{+}$. Replacing $x\in X_{+}$ with $nx$ for large $n$ we get that $0\leq\langle J^{\prime}z^{\prime},x\rangle$, so $J^{\prime}z^{\prime}\geq 0$, contradicting the bipositivity of $J^{\prime}$.
Conversely, let $J(X_{+})$ be dense in $Z_{+}$. In particular $J$ and hence, $J^{\prime}$ is positive. Now let $z^{\prime}\in Z^{\prime}$ be such that $J^{\prime}z^{\prime}\geq 0$. Therefore, $\langle z^{\prime},Jx\rangle\geq 0$ for all $x\in X_{+}$. By density of $J(X_{+})$ in $Z_{+}$, we conclude that $\langle z^{\prime},z\rangle\geq 0$ for all $z\in Z_{+}$, so $z^{\prime}\geq 0$.
(c) Note that $J^{\prime}(Z^{\prime}_{+})\subseteq X^{\prime}_{+}$, so we only need to show the reverse inclusion. For this, let $x^{\prime}\in X^{\prime}_{+}$. The linear mapping $\varphi:J(X)\to\mathbb{R}$, $v\mapsto\langle x^{\prime},J^{-1}v\rangle$ is well-defined and positive by the injectivity and bipositivity of $J$. Since $J(X)$ is by assumption majorizing in $Z$, the Kantorovich extension theorem [4, Theorem 1.36] (which is formulated there for ordered vector spaces, but can be checked to also hold in the pre-ordered case) implies that $\varphi$ extends to a positive linear functional $z^{\prime}:Z\to\mathbb{R}$. As $Z_{+}$ is generating, $z^{\prime}$ is automatically continuous [7, Theorem 2.8], i.e., $z^{\prime}\in Z^{\prime}_{+}$. Moreover, the equality $\langle z^{\prime},Jx\rangle=\varphi(Jx)=\langle x^{\prime},x\rangle$ for each $x\in X$ implies that $x^{\prime}=J^{\prime}z^{\prime}\in J^{\prime}(Z^{\prime}_{+})$. ∎
@@
For $k\in\mathbb{N}$ and a bounded domain $\Omega\subseteq\mathbb{R}^{d}$, let $\mathrm{C}^{k}(\overline{\Omega})$ denote the space of all $k$-times continuously differentiable functions on $\Omega$ whose derivatives up to order $k$ extend continuously to $\overline{\Omega}$. This is an ordered Banach space with the usual pointwise order and norm. We show that its dual cone is (more precisely, can be identified with) the set of all positive finite Borel measures on $\overline{\Omega}$.
Indeed, the canonical embedding $J:\mathrm{C}^{k}(\overline{\Omega})\to\mathrm{C}(\overline{\Omega})$ is bipositive and has majorizing range because it contains the constant function $\operatorname{{\mathbbm{1}}}$. Moreover, by employing the density of $\mathrm{C}^{k}(\overline{\Omega})$ in $\mathrm{C}(\overline{\Omega})$ (which is true by the Stone-Weierstraß approximation theorem) and by using shifts by small multiples of the constant function $\operatorname{{\mathbbm{1}}}$, we even have that the cone $\mathrm{C}^{k}(\overline{\Omega})_{+}$ is dense in the cone $\mathrm{C}(\overline{\Omega})_{+}$. So $J^{\prime}$ is bipositive and injective according to Proposition 2.1. Hence, $J^{\prime}$ is a bipositive embedding of $\mathrm{C}(\overline{\Omega})^{\prime}$ (which can be identified with the space of finite Borel measures on $\overline{\Omega}$) into $\big{(}\mathrm{C}^{k}(\overline{\Omega})\big{)}^{\prime}$. Hence, Proposition 2.1(c) shows that the dual cone of $\mathrm{C}^{k}(\overline{\Omega})$ consists of (to be precise, can be identified via $J^{\prime}$ with) the set of all positive finite Borel measures on $\overline{\Omega}$.
In our next result, we show that the first two properties in Table 1 are preserved by the inverse of a positive bounded bijection (which might not be bipositive).
@@
    If the cone of $X$ is a face of its bidual wedge, then the cone of $Z$ is also a face of its bidual wedge.
@@
Proof.
(a) Let $(z_{\alpha})$ be an increasing norm-bounded net in $Z_{+}$. Then $(Tz_{\alpha})$ is increasing and norm-bounded in $X_{+}$. So, by assumption, there exists $x\in X$ such that $Tz_{\alpha}\to x$ in $X$. The bounded inverse theorem now implies that $z_{\alpha}\to T^{-1}x$ in $Z$, as desired.
(b) Let $k:X\to X^{\prime\prime}$ and $\widetilde{k}:Z\to Z^{\prime\prime}$ be the canonical embeddings. Let $z\in Z_{+}$ and $z^{\prime\prime}\in Z^{\prime\prime}_{+}$ such that $0\leq z^{\prime\prime}\leq\widetilde{k}(z)$. We need to show that $z^{\prime\prime}\in\widetilde{k}(Z)$. Since $T$ is positive, so is the double dual $T^{\prime\prime}:Z^{\prime\prime}\to X^{\prime\prime}$. Therefore, $0\leq T^{\prime\prime}z^{\prime\prime}\leq T^{\prime\prime}\widetilde{k}(z)=k(Tz)$. As $k(X_{+})$ is a face of $X^{\prime\prime}_{+}$, it follows that $T^{\prime\prime}z^{\prime\prime}\in k(X)$. Hence, $z^{\prime\prime}\in(T^{\prime\prime})^{-1}k(X)=\widetilde{k}(T^{-1}X)=\widetilde{k}(Z)$.
(c) First of all, recall that normality of the cone in an ordered Banach space is equivalent to every order interval being norm-bounded [4, Theorem 2.40]. Now, let $a,b\in Z$. Then $T[a,b]\subseteq[Ta,Tb]$ due to the positivity of $T$ and the later set is norm-bounded because of the normality of the cone in $X$. Hence, $[a,b]=T^{-1}T[a,b]$ is also norm-bounded since $T^{-1}$ is continuous. ∎
Regarding part (b) of the previous theorem, we note in passing that if a cone is a face in a wedge, then it follows that the latter is also a cone.
For a bipositive operator $J:X\to Z$ between ordered Banach spaces, we give conditions in Theorem 2.5 that ensure that $J(X_{+})$ is a face of $Z_{+}$. For the proof we need the following extension result for linear functionals.
@@
Let $X$ be a pre-ordered Banach space with a generating wedge and let $V\subseteq X$ be a vector subspace. Let $x^{\prime}\in X^{\prime}_{+}$ and $\varphi:V\to\mathbb{R}$ be linear. The following are equivalent.
@@
“(ii) $\Rightarrow$ (i)”: Firstly, note that the functional $p:X\to[0,\infty)$ given by $x\mapsto\inf\{\left\langle x^{\prime}\,,\,w\right\rangle:w\in X_{+},w\geq x\}$ is well-defined since $X_{+}$ is generating and can easily checked to be sublinear. Moreover, by (ii) one has $\varphi(v)\leq p(v)$ for all $v\in V$. The Hahn-Banach theorem thus yields a linear functional $y^{\prime}:X\to\mathbb{R}$ that satisfies $y^{\prime}\raisebox{-2.15277pt}{$|$}_{V}=\varphi$ and $y^{\prime}(x)\leq p(x)$ for all $x\in X$. For each $x\in X_{+}$, observe that $y^{\prime}(-x)\leq p(-x)=0,$ so $y^{\prime}$ is positive. In particular, $y^{\prime}\in X^{\prime}$ since $X_{+}$ is generating [7, Theorem 2.8]. Finally, $\left\langle y^{\prime}\,,\,x\right\rangle\leq p(x)\leq\left\langle x^{\prime}\,,\,x\right\rangle$ for all $x\in X_{+}$ proves that $y^{\prime}\leq x^{\prime}$. ∎
@@
Let $X$ be an ordered Banach space with normal cone, let $Z$ be a pre-ordered Banach space, and let $J\in\mathcal{L}(X,Z)_{+}$ have dense range. Assume that $(R_{n})$ is a sequence in $\mathcal{L}(Z,X)_{+}$ such that $(JR_{n})$ and $(R_{n}J)$ both converge to the identity operator in the weak operator topology on $\mathcal{L}(Z)$ and $\mathcal{L}(X)$ respectively.
If $X_{+}$ is a face of $X^{\prime\prime}_{+}$, then $J(X_{+})$ is a face of $Z_{+}$ (in particular, $Z_{+}$ is a cone).
We point out that the map $J$ in the theorem is automatically bipositive: if $Jx\geq 0$, then $x$ is the weak limit of $(R_{n}Jx)\subseteq X_{+}$, hence positive. Moreover, Example 2.11 shows that the assumption that $X_{+}$ is a face of $X^{\prime\prime}_{+}$ cannot be dropped in Theorem 2.5.
Proof of Theorem 2.5.
We denote by $k:X\to X^{\prime\prime}$ and $\widetilde{k}:Z\to Z^{\prime\prime}$ the canonical embeddings. Then the following diagram commutes:
@@
Let $x\in X_{+}$ and $z\in Z_{+}$ be such that $0\leq z\leq Jx$; we need to find $y\in X_{+}$ such that $Jy=z$. To this end, we first find a $y^{\prime\prime}\in X^{\prime\prime}_{+}$ such that $J^{\prime\prime}y^{\prime\prime}=\widetilde{k}(z)$.
For the first step, define $\varphi:J^{\prime}(Z^{\prime})\to\mathbb{R}$ as $v^{\prime}\mapsto\langle\widetilde{k}(z),(J^{\prime})^{-1}v^{\prime}\rangle.$ This is well-defined since $J^{\prime}$ is injective as $J$ is assumed to have dense range. We want to extend $\varphi$ to a functional $y^{\prime\prime}\in X^{\prime\prime}_{+}$ by employing Proposition 2.4. So let $v^{\prime}=J^{\prime}z^{\prime}\in J^{\prime}(Z^{\prime})\subseteq X^{\prime}$ and fix $w^{\prime}\in X^{\prime}_{+}$ such that $w^{\prime}\geq v^{\prime}$. Then
$\varphi(v^{\prime})=\left\langle\widetilde{k}(z)\,,\,z^{\prime}\right\rangle=\left\langle z^{\prime}\,,\,z\right\rangle=\lim_{n\to\infty}\left\langle z^{\prime}\,,\,JR_{n}z\right\rangle=\lim_{n\to\infty}\left\langle v^{\prime}\,,\,R_{n}z\right\rangle.$
Positivity of $z$ and $w^{\prime}$ thus implies that
$\varphi(v^{\prime})\leq\limsup_{n\to\infty}\left\langle w^{\prime}\,,\,R_{n}z\right\rangle\leq\limsup_{n\to\infty}\left\langle w^{\prime}\,,\,R_{n}Jx\right\rangle=\left\langle w^{\prime}\,,\,x\right\rangle=\left\langle k(x)\,,\,w^{\prime}\right\rangle.$
Thus by Proposition 2.4 – which is applicable since $X^{\prime}_{+}$ is generating as $X_{+}$ is normal [4, Theorem 2.40] – there exists $y^{\prime\prime}\in X^{\prime\prime}$ such that $0\leq y^{\prime\prime}\leq k(x)$ and $y^{\prime\prime}\raisebox{-2.15277pt}{$|$}_{J^{\prime}(Z^{\prime})}=\varphi$. Now, $\widetilde{k}(z)=J^{\prime\prime}y^{\prime\prime}$, because for each $z^{\prime}\in Z^{\prime}$, we have
$\left\langle\widetilde{k}(z)\,,\,z^{\prime}\right\rangle=\varphi(J^{\prime}z^{\prime})=\left\langle y^{\prime\prime}\,,\,J^{\prime}z^{\prime}\right\rangle=\left\langle J^{\prime\prime}y^{\prime\prime}\,,\,z^{\prime}\right\rangle.$
Now, we show that $y^{\prime\prime}$ even stems from an element of $X$. Since $k(X_{+})$ is a face of $X^{\prime\prime}_{+}$ and $0\leq y^{\prime\prime}\leq k(x)$, there exists $y\in X$ such that $y^{\prime\prime}=k(y)$. As $k$ is bipositive, it follows that $0\leq y\leq x$. Moreover, as $\widetilde{k}(z)=J^{\prime\prime}k(y)=\widetilde{k}(Jy)$, we conclude $z=Jy$. ∎
If $X$ and $D$ are Banach spaces, then we know as a consequence of Hahn-Banach theorem that every $T\in\mathcal{L}(X,D)$ satisfies the norm equality $\left\lVert T^{\prime}\right\rVert_{D^{\prime}\to X^{\prime}}=\left\lVert T\right\rVert_{X\to D}$. The following theorem can be interpreted as a variation of the inequality $\leq$” in this equality. For an intuition about the theorem, we refer to the subsequent Examples 2.7.
@@
for a number $c\geq 0$ that might depend on all involved spaces and operators except for $X$ and $T$. Here, $\operatorname{span}E^{\prime}_{+}=E^{\prime}_{+}-E^{\prime}_{+}$ is endowed with the norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{E^{\prime}_{+}-E^{\prime}_{+}}$.
Note that the space $D$ does not occur in the conclusion of the theorem. We are interested in the mappings $\tilde{J}T:X\to\tilde{E}$ and $(JT)^{\prime}:E^{\prime}\to X^{\prime}$ and $D$ is an auxiliary space to relate their properties. Before giving the proof of the theorem, we illustrate two situations where the assumptions of the theorem are fulfilled:
Examples 2.7.
(a) Let $p\in[1,\infty]$, let $T\in\mathcal{L}(L^{p}(\mathbb{R}),L^{\infty}(\mathbb{R}))_{+}$, and assume that the range of $T$ lies in $\operatorname{Lip}(\mathbb{R})$, the space of all scalar-valued Lipschitz continuous functions on $\mathbb{R}$, which is an ordered Banach space when endowed with the pointwise order and the norm $\left\lVert f\right\rVert_{\operatorname{Lip}}:=\left\lvert f(0)\right\rvert+\operatorname{Lip}(f)$, where $\operatorname{Lip}(f)$ is the Lipschitz constant of $f$. Then Theorem 2.6 implies that there exists $c>0$, independent of $T$ such that
$\left\lVert T^{\prime}\right\rVert_{\operatorname{span}(\operatorname{Lip}(\mathbb{R})^{\prime}_{+})\to(L^{p})^{\prime}}\leq c\left\lVert T\right\rVert_{L^{p}\to L^{\infty}}.$
To see this, let $X=L^{p}(\mathbb{R})$, $E=\operatorname{Lip}(\mathbb{R})$, $\tilde{E}=L^{\infty}(\mathbb{R})$ and $D=E\cap\tilde{E}$ with the norm $\left\lVert f\right\rVert_{D}:=\left\lVert f\right\rVert_{\operatorname{Lip}}+\left\lVert f\right\rVert_{\infty}$ in the theorem, let $J$ and $\tilde{J}$ be the canonical embeddings, and note that $T$ is bounded from $L^{p}(\mathbb{R})$ to $D$ by the closed graph theorem. Moreover, $D$ is majorizing in $\tilde{E}$ as the former contains the constant $\operatorname{{\mathbbm{1}}}$ function.